OpenStudy (anonymous):

-1 3/7 times -3 2/3 I got: 3 2/7. Am I correct? Will medal and fan!

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

@sammixboo @inkyvoyd

OpenStudy (anonymous):

Um, no. Your answer is incorrect. So, I made those mixed fractions into improper fractions. -1 3/7 = -10/7 -3 2/3 = -11/3 Multiply -10/7 and -11/3. = 110/21 Reduce 110/21. = 5 5/21 The correct answer is 5 5/21.

OpenStudy (anonymous):

Oh wow. Thanks. I just have one more.